# Product Digital Twin Supporting End-of-life Phase of Electric Vehicle Batteries Utilizing Product-Process-Resource Asset Network

### v1

@@ Abstract
In the context of the circular economy, products in their end-of-life phase should be either remanufactured or recycled. Both of these processes are crucial for sustainability and environmental conservation. However, manufacturers often do not support these processes enough by not sharing relevant data. This paper proposes use of a digital twin technology, which is capable to help optimizing the disassembly processes to reduce ecological impact and enhance sustainability. The proposed approach is demonstrated through a disassembly use-case of the product digital twin of an electric vehicle battery. By utilizing product digital twins, challenges associated with the disassembly of electric vehicle batteries can be solved flexibly and efficiently for various battery types. As a backbone for the product digital twin representation, the paper uses the paradigm of product-process-resource asset networks (PAN). Such networks enable to model relevant relationships across products, production resources, manufacturing processes, and specific production operations that have to be done in the manufacturing phase of a product. This paper introduces a Bi-Flow Product-Process-Resource Asset Network (Bi-PAN) representation, which extends the PAN paradigm to cover not only the manufacturing, but also the remanufacturing/recycling phase.
@@ I Introduction
The circular economy is an economic system that aims to keep resources in use for as long as possible, through a closed-loop approach of reducing, reusing, remanufacturing, and recycling materials and products. It contrasts with the traditional linear economy, where resources are extracted, transformed into products, used, and disposed of as waste [1].
@@
Recycling electric vehicle (EV) batteries is crucial for the circular economy because it can help to reduce waste and conserve resources. EV batteries contain valuable materials such as lithium, cobalt, nickel or others depending on the specific battery chemistry, which can be recovered and reused. By recycling these materials, the need for mining new resources is reduced, which can have positive environmental and social impacts [2]. Conventional methods for EV battery recycling are based on crushing or shredding the entire battery or battery modules, separating the chemical elements and their reusing for new batteries. On the contrary, another approach is based on reusing battery modules in a satisfactory state-of-health for stationary applications in households, industrial factories, energy sources, or energy storages [3]. This is the case that motivates the presented research. To be able to get battery modules in a good state, a support for dismantling the entire EV battery is needed. Moreover, recycling or reusing EV batteries can also provide a solution for managing their end-of-life. As the number of EVs on the road is continually increasing, the number of batteries reaching their end-of-life will also increase. If these batteries are not properly managed, they can pose environmental and safety risks.
@@
Production processes can be efficiently supported by a technology of digital twins, which pose digital replicas for physical systems. In the domain of EV batteries, integration with digital twins is already a researched matter. Articles [4] and [5] provide extensive information on digital twin for EV batteries in several possible applications and usages as well as highlighting the advantages of using digital twins for remanufacturing and recycling EV batteries.
This approach highlights the use of digital twins in the end-of-life phase of EV batteries utilizing the so-called Product Digital Twin (PDT). PDT is enhancing the sustainability and efficiency of product lifecycle management processes. PDT brings to manufacturers and remanufacturers a comprehensive tool for planning and optimizing products.
@@
In this paper, the concept of a Bi-Flow Product-Process-Resource Asset Network (Bi-PAN) is introduced. Bi-PAN is designed to support both assembly and disassembly workflows within a single framework. Built on the foundation of the Product-Process-Resource Asset Network (PAN) [6], the extended framework preserves the representation of products, processes, and resources accompanied by their links, while incorporating also additional features to support reverse process flows. Thus, Bi-PAN supports the end-of-life phase of a product seamlessly together with production data. By integrating disassembly capabilities into the PAN, it serves to efficiently facilitate the design of remanufacturing and recycling processes in addition to the assembly process.
The Bi-PAN and Product Digital Twin together offer a holistic approach to product lifecycle optimization, facilitating the transition towards more sustainable and circular manufacturing practices.
@@ II State of the Art
The proposed approach is built on the top of 2 main research areas including (i) digital twin and Product Life-Cycle, and (ii) Product-Process-Resource Asset Network. These domains are described in details in the following subsections.
@@ II-A Digital Twin and Product Life-Cycle
Digital Twin is a technology that has widespread in various industries and applications [7]. It is a digital replica or virtual representation of a physical object, system, or process. A digital twin can dynamically mirror and capture the characteristics of its real-world counterpart in real time. It can also provide ways to simulate, analyze, maintain, and optimize the physical entity’s operation and behavior in virtual environment [8]. Moreover, it is important to note that defining a digital twin unambiguously can be challenging. Reference [9] provides a collection of various characterizations of digital twin, offering insights into the diverse interpretations.
Due to the fact that digital twin enables the storage and management of information and data about the products, including their design, materials, components, manufacturing processes, and maintenance history, digital twin is capable of product life-cycle management throughout the entire lifespan of the products Fig. 1. The key feature of digital twin for its use in the product life cycle management is its nature as continuously evolving entity, capable of providing ongoing value throughout its lifespan. They can be stored indefinitely and potentially offer valuable insights for future analysis, even after the physical entity has been disposed of [10].
Figure 1: Visualisation of the product digital twin life cycle captures the journey from the initial product design to the end-of-life, closing the loop by offering insights gathered throughout the product life cycle back into the design process.
Digital twin offers diverse applications, leading to the development of various frameworks and methodologies. These frameworks encompass a range of technologies, reflecting the versatility of digital twin creation. In the context of product life-cycle management, the framework proposed in reference [11] highlights the significance of an Asset Administration Shell (AAS) [12, 13] as an option that can provide sustainable solutions choice while also ensuring interoperability. AAS, being a framework, requires an actual technology for digital twin creation within its parameters. AutomationML [14], for instance, offers an XML-based language tailored for describing automation systems. Utilizing AutomationML within the AAS framework [15] facilitates the creation of digital twins while adhering to its principles.
@@ II-B Product-Process-Resource Asset Network
The Product-Process-Resource Asset Network (PAN) [6] is a graph-based formalism for expressing relationships among products, production processes, and resources. The PAN modeling is suitable for capturing and sharing knowledge among engineers and respective engineering tools as well as for identification of changes during engineering processes of multi-disciplinary system engineering projects. In fact, the PAN approach revolutionizes engineering processes by virtualizing assets, operations, and resources and especially by mapping and connecting them together by explicitly specified links. Usage of PAN optimizes efficiency of engineering processes, it can support sustainability, and it enhances operations by modeling manufacturing processes, tracking resource usage, and monitoring asset performance. Through data integration, visualization, and analysis, PAN enables informed decision-making [16]. The PAN concept builds upon the Product-Process-Resource (PPR) [17] concept/categorization, which consolidates product, process, and resource information into Engineering Assets. However, in the domain of Cyber-Physical Production Systems (CPPSs), these dependencies are often left implicit. The PAN addresses this gap by providing a structured framework for explicitly articulating and overseeing the dependencies among product, process, and resource assets. We are not aware of using PANs as a core backbone for designing digital twins for products or processes in flexible manufacturing or remanufacturing, as we do in this paper.
@@ III Product Digital Twin and Bi-Flow Product-Process-Resource Asset Network
This paper addresses the concept of the digital twin for a product, hereinafter called the Product Digital Twin (PDT). The PDT represents a digital twin specifically designed for life-cycle management of products. It facilitates easy access to relevant product information, including data from manufacturers such as maintenance and service details, operational data or mechanical composition, as well as data gathered throughout the product’s lifecycle. This paper highlights the significance of information relevant for the end-of-life phase of products, as it represents a crucial opportunity to achieve sustainability. Specifically, the scope is focused on recycling and remanufacturing of products - both in terms of repair and re-commissioning, and in terms of transformation into a new product. By leveraging data from manufacturers and accumulated lifecycle data, PDT aims to enhance the efficiency and sustainability of remanufacturing and recycling processes.
The benefits of the usage of the PDT in the end-of-life phase includes:
@@
    Designing remanufacturing/recycling process: PDT enables the design, planning and simulation of the remanufacturing/recycling process before the physical product arrives. This pre-planning can significantly improve the efficiency of the remanufacturing/recycling process.
@@
    Individualized approach: Each product having its own PDT allows for an individualized approach during remanugacturing/recycling process. This is crucial because products may undergo changes or experience defects during their life-cycle. The PDT ensures that each product is handled appropriately based on its specific characteristics and condition.
@@
    Real-time monitoring and adaptation: PDT provides real-time monitoring of the product’s status during remanufacturing/recycling process. This capability is valuable for quick dealing with unexpected events or errors in the process.
@@
    Data collection: The data collected during the remanufacturing and recycling process can be invaluable for manufacturers. They can analyze this data to gain insights into how products degrade or wear over time or to analyze problems during the remanufacturing/recycling process. This information can inform the design of the next generation of products, making them more sustainable. For remanufacturers/recyclers this data could be used to make remanufacturing/recycling process more efficient.
While the advantages and applications of PDT are becoming increasingly clear, a significant challenge lies in creating a PDT that can effectively carry all the relevant information. When creating a PDT, it’s crucial to clearly define the dependencies among individual components. This includes determining the sequence of component removal, understanding the method of removal (e.g., bolts, clips), and considering factors like accessibility and interconnectivity. These details are essential for automating the disassembly process. The PDT should be structured to represent the product’s components and their relationships, rather than detailing specific disassembly procedures, as multiple disassembly strategies may exist. To address this need, the content of the Product Digital Twin (PDT) includes a generic disassembly plan. To visualize such a plan, an Bi-Flow Product-Process-Resource Asset Network (Bi-PAN) framework is employed. This Bi-PAN framework not only encompasses links between products and processes but also considers the resources required for each process. It facilitates the visualization of both assembly and disassembly processes, providing insights into the steps involved and the resources needed for efficient disassembly of the product.
Figure 2: Bi-PAN for the BMW i3 EV battery, representing product stages, production process operations and resource skills. The legend explaining the nomenclature is included in the bottom part of the figure.
Just like the classic PAN, the Bi-PAN contains the same three basic elements (see Fig. 2): products as red circles, processes as green rectangles, and resources as blue rounded rectangles, all interconnected to represent their relationships. Black arrows show the connection between individual products and processes, forming an assembly or disassembly plan. Yellow markers link process steps with resources or resource skills, assigning specific assembly or disassembly procedure to individual processes. Depending on the direction of movement, directional arrows appropriately indicate the flow. Within the Bi-PAN, products are categorized based on their assembly or disassembly role as follows:
@@ IV Product Digital Twin Based on the Bi-PAN for the Electric Vehicle Battery
An EV battery serves as an exemplary case for demonstrating the PDT based on the Bi-PAN approach. Throughout the EV battery lifecycle, all three scenarios of the end-of-life phase are relevant: recycling to reintegrate scarce materials within the circular economy, repairing or replacing non-functional parts for re-use of EV batteries, and transforming EV batteries into modular units for secondary energy-storage applications. Remanufacturing or recycling of EV batteries also presents challenges, including the presence of hazardous substances and the risk of spontaneous combustion. It is therefore important to have relevant information and to select the appropriate remanufacturing/recycling route accurately and efficiently. This underscores the importance of PDT.
@@
When it comes to selecting the appropriate recycling process for an EV battery, understanding the battery structure is paramount. This entails not just knowing the list of individual components but also comprehending the interdependencies between these components, including their composition. For the creation of the PDT, the utilization of the AutomationML tool proves invaluable. This tool enables not only the listing of individual components of the EV battery but also the depiction of their dependencies and detailed information. For instance, it can provide data about each component’s position in the battery coordinate system, which proves beneficial when considering disassembly by a robot. Example of PDT and its structure in AutomationML system unit class is depicted in Fig. 3. The data obtained from AutomationML is important, but for the ability to plan assembly or disassembly operations we lack the linking of the already described product structure with specific processes and resources. The Bi-PAN will serve us for this purpose.
@@
Figure 3: Example of system unit class in AutomationML as a screenshot from the AutomationML Editor.
@@
As already introduced in Sec. III, Bi-PAN includes connections between products and processes, while also assigns the resources needed for each process. The Bi-PAN for the EV battery is depicted in the already mentioned Fig. 2. The following description explains the assembly process of an EV battery that is starting from the top, and therefore the relevant direction for the black arrows is downwards. The assembly commences with three separate products: screws (Fastener Product), the cooling system (SubProduct), and the battery box (Elementary Product). These products are inserted into a single assembly through a process requiring resource skills for manipulation and screwing, as indicated by yellow markers. The output of this process is Stage 1 (i.e., Product in the Stage in the Bi-PAN notation) of the EV battery assembly. The Stage 1 then proceeds to the next process along with bolts, BMS & cables, and eight modules, where they are once again inserted into a single assembly. This process involves resource skills manipulation, screwing, and connecting cables. The result is Stage 2 of the EV battery assembly. Stage 2 and bolts, along with five braces, proceed to the next process where they are inserted into a single unit. This process requires resource skills for manipulation and screwing. The output of this process is Stage 3 of the EV battery assembly process. Stage 3, along with the isolation blanket, enters the subsequent process, requiring resource skills for manipulation only. This creates Stage 4 of the EV battery assembly. Finally, Stage 4, along with screws and a cover, enters the last process, requiring resource skills for manipulation and screwing. The output of this process is the Final Product - the assembled EV battery.
@@
To illustrate the application of the Bi-PAN in remanufacturing, we’ve selected the scenario of replacing a broken module with a new one. This specific process represents a sequence of operations that commences in a particular state of the EV battery and concludes at the desired state. The remanufacturing process depicted in Fig. 4 is derived from the generic Bi-PAN for the entire EV battery (i.e., from Bi-PAN in Fig. 2, which is repeated without labels on the left-hand side of the figure). The flow of the whole remanufacturing process is depicted by red arrow starting from the EV battery (Final Product) continuing as a disassembly process to the Stage 2, where the damaged module is removed and replaced by the new one and the it continues as an assembly process and ends as an EV battery (Final Product).

### v2

# Product Digital Twin Supporting End-of-life Phase of Electric Vehicle Batteries Utilizing Product–Process–Resource Asset Network

@@ Abstract
In a circular economy, products in their end-of-life phase should be either remanufactured or recycled. Both of these processes are crucial for sustainability and environmental conservation. However, manufacturers frequently do not support these processes enough in terms of not sharing relevant data about the products nor their (re-)manufacturing processes. This paper proposes to accompany each product with a digital twin technology, specifically the Product Digital Twin (PDT), which can carry information for facilitating and optimizing production and remanufacturing processes. This paper introduces a knowledge representation called Bi-Flow Product-Process-Resource Asset Network (Bi-PAN). Bi-PAN extends a well-proven Product-Process-Resource Asset Network (PAN) paradigm by integrating both assembly and disassembly workflows into a single information model. Such networks enable capturing relevant relationships across products, production resources, manufacturing processes, and specific production operations that have to be done in the manufacturing phase of a product. The proposed approach is demonstrated in a use-case of disassembling electric vehicle (EV) batteries. By utilizing PDTs with Bi-PAN knowledge models, challenges associated with disassembling of EV batteries can be solved flexibly and efficiently for various battery types, enhancing the sustainability of the EV battery life-cycle management.
@@ I Introduction
A circular economy is an economic system aiming at retaining resources in usage for as long as possible, realizing a closed-loop approach of reducing, reusing, remanufacturing, and recycling materials and products. It contrasts with the traditional linear economy, where resources are extracted, transformed into products, utilized, and thrown away as waste [1]. Recycling electric vehicle (EV) batteries is crucial for the circular economy because it can help reducing waste and conserve resources. EV batteries contain valuable materials such as lithium, nickel, cobalt or others depending on the specific battery chemistry, which can be recovered and reused. By recycling these materials, the need for mining new resources is reduced, which can have positive environmental and social impacts [2]. Conventional methods for EV battery recycling are based on crushing or shredding the entire battery or battery modules, separating the chemical elements, and reusing them for new batteries. On the contrary, another approach is based on reusing battery modules in a satisfactory state-of-health for stationary applications in households, industrial factories, energy sources, or energy storages [3]. This is the case that motivates the presented research. To be able to get battery modules in a good state, support for dismantling the entire EV battery is needed. Moreover, recycling or reusing EV batteries can also provide a solution for managing their end-of-life. As the count of EVs driving on the road is continually increasing, the number of batteries reaching their end-of-life will also increase. If these batteries are not properly managed, they can pose environmental and safety risks.
@@
Production processes can be efficiently supported by a technology of digital twins, which pose digital replicas for physical systems [4, 5]. In the domain of EV batteries, integration with digital twins is already a researched matter. Articles [6] and [7] provide extensive information on digital twins for EV batteries in several possible applications and usages as well as highlighting the benefits of using digital twins for remanufacturing and recycling EV batteries. This paper focuses on the application of digital twins in the end-of-life phase of EV batteries utilizing the so-called “Product Digital Twin” (PDT). PDTs improve sustainability and efficiency in product life-cycle [8] management by offering manufacturers and remanufacturers a comprehensive tool for planning and optimizing the end-of-life processes of these products.
@@
For representing PDTs, this paper proposes a concept of a Bi-Flow Product-Process-Resource Asset Network (Bi-PAN), building on the Product-Process-Resource Asset Network (PAN) [9]. PAN is a model that integrates product description, production processes, and resource information into a cohesive network/graph, modeling relevant relationships among these entities to support efficient planning, coordination, and optimization throughout production [10]. PAN is designed in compliance with the industrial standard VDI/VDE 3682, which is widely accepted in the industry, especially in the automotive sector. Bi-PAN extends PAN by supporting both assembly and disassembly workflows within a single model, preserving the representation of products, processes, and resources along with their interconnections, while also incorporating additional features to handle reverse process flows. Thus, Bi-PAN supports the end-of-life phase of a product seamlessly together with production data. By integrating disassembly capabilities into the PAN, it serves to efficiently facilitate the design of remanufacturing and recycling processes in addition to the assembly process. The Product Digital Twin using the Bi-PAN model offers a holistic approach to product life-cycle optimization, facilitating the transition towards more sustainable and circular manufacturing practices.
@@ II State of the Art
The proposed approach is built on the top of two main research areas including (i) Digital Twin and Product Life-Cycle, and (ii) Product-Process-Resource Asset Network. These domains are described in detail in the following subsections.
@@ II-A Digital Twin and Product Life-Cycle
Digital Twin technology has widespread use in various industries and applications [11]. It is a digital replica or virtual representation of a physical object, system, or process. A digital twin can dynamically mirror and capture the characteristics of its real-world counterpart in real-time. It can also provide ways to simulate, analyze, maintain, and optimize the physical entity’s operation and behavior in a virtual environment [12]. Moreover, defining a digital twin unambiguously can be challenging. Reference [13] provides a collection of various characterizations of digital twins, offering insights into the diverse interpretations.
Due to the fact that digital twin enables the storage and management of information and data about the products, including their design, materials, components, manufacturing processes, and maintenance history, digital twin is capable of product life-cycle management throughout the entire lifespan of the products. The key feature of a digital twin for its use in product life-cycle management is its nature as a continuously evolving entity, capable of providing ongoing value throughout its lifespan. They can be stored indefinitely and potentially offer valuable insights for future analysis, even after the physical entity has been disposed of [14].
Digital twin offers diverse applications, leading to the development of various frameworks and methodologies. These frameworks encompass a range of technologies, reflecting the versatility of digital twin creation. In the context of product life-cycle management, the approach proposed in reference [15] highlights the significance of an Asset Administration Shell (AAS) [16, 17] as an option that can provide sustainable solutions choice while also ensuring interoperability. AAS, being a framework, requires an actual technology for digital twin creation within its parameters. AutomationML [18], for instance, offers an XML-based language tailored for describing automation systems. Utilizing AutomationML within the AAS framework [19] facilitates the creation of digital twins while adhering to its principles.
@@ II-B Product-Process-Resource Asset Network
The Product-Process-Resource Asset Network (PAN) [9] is a graph-based formalism for expressing relationships among products, production processes, and resources. The PAN modeling is suitable for capturing and sharing knowledge among engineers and respective engineering tools as well as for identification of changes during engineering processes of multi-disciplinary system engineering projects. In fact, the PAN approach revolutionizes engineering processes by virtualizing assets, operations, and resources and especially by mapping and connecting them together by explicitly specified links. The usage of PAN optimizes the efficiency of engineering processes, it can support sustainability and enhances operations by modeling manufacturing processes, tracking resource usage, and monitoring asset performance. Through data integration, visualization, and analysis, PAN enables informed decision-making [10]. The PAN concept builds upon the Product-Process-Resource (PPR) [20] concept/categorization, which consolidates product, process, and resource information into Engineering Assets. However, in the domain of Cyber-Physical Production Systems (CPPSs), these dependencies are frequently left implicit. The PAN addresses this gap by providing a structured model for explicitly articulating and overseeing the dependencies among product, process, and resource assets. We have been investigating possible extensions of the PAN paradigm by justifying the focus on a product structure, which has been addressed in [21].
@@ III Product Digital Twin and Bi-Flow Product-Process-Resource Asset Network
This paper addresses the concept of the digital twin for a product, hereinafter called the Product Digital Twin (PDT). The PDT represents a digital twin specifically designed for life-cycle management of products. It facilitates easy access to relevant product information, including data from manufacturers such as maintenance and service details, operational data, or mechanical composition, as well as data gathered throughout the product’s life-cycle. This paper highlights the significance of information relevant for the end-of-life phase of products, as it represents a crucial opportunity to achieve sustainability. Specifically, the scope is focused on recycling and remanufacturing of products - both in terms of repair and re-commissioning and in terms of transformation into a new product. By leveraging data from manufacturers and accumulated life-cycle data, PDT aims at strengthening the efficiency and sustainability of remanufacturing and recycling processes.
The benefits of the usage of the PDT in the end-of-life phase include:
@@
    Designing remanufacturing/recycling process: PDT enables the design, planning, and simulation of the remanufacturing/recycling process before the physical product arrives. This pre-planning can significantly improve the efficiency of the remanufacturing/recycling process.
@@
    Individualized approach: Each product having its own PDT allows for an individualized approach during the remanufacturing/recycling process. This is crucial because products may undergo changes or experience defects during their life-cycle. The PDT ensures that each product is handled appropriately based on its specific characteristic and condition.
@@
    Real-time monitoring and adaptation: PDT provides online monitoring of the product’s status during remanufacturing/recycling process. This capability is valuable for quick dealing with unexpected events or errors in the process.
@@
    Data collection: The data collected during the remanufacturing and recycling process can be invaluable for manufacturers. They can analyze this data to gain insights into how products degrade or wear over time or to analyze problems during the remanufacturing/recycling process. This information can inform the design of the next generation of products, making them more sustainable. For remanufacturers/recyclers, this data could be used to make the remanufacturing/recycling process more efficient.
While the advantages and applications of PDT are becoming increasingly clear, a significant challenge lies in creating a PDT that can effectively carry all the relevant information. When creating a PDT, it’s crucial to clearly define the dependencies among individual components. This includes determining the sequence of component removal, understanding the method of removal (e.g., bolts, clips), and considering factors like accessibility and interconnectivity. These details are essential for automating the disassembly process. The PDT should be structured to represent the product’s components and their relationships, rather than detailing specific disassembly procedures, as multiple disassembly strategies may exist. To address this need, the content of the Product Digital Twin (PDT) includes a generic disassembly plan. To visualize such a plan, a Bi-Flow Product-Process-Resource Asset Network (Bi-PAN) model is employed. This Bi-PAN model not only encompasses links between products and processes but also considers the resources required for each process. It facilitates the visualization of both assembly and disassembly processes, providing insights into the steps involved and the resources needed for efficient disassembly of the product.
Figure 1: Bi-PAN for the simplified BMW i3 EV battery, representing product stages, production process operations, and resource skills. The legend explaining the nomenclature is included in the bottom part of the figure.
Just like the classic PAN, the Bi-PAN contains the same three basic elements (see Fig. 1): products as red circles, processes as green rectangles, and resources as blue rounded rectangles, all interconnected to represent their relationships. Black arrows show the connection between individual products and processes, forming an assembly or disassembly plan depending on the shape of the arrow. Open arrows represent assembly direction, while full arrows represent disassembly direction. Yellow markers link process steps with resources or resource skills, assigning specific assembly or disassembly procedures to individual processes. Depending on the direction of movement, directional arrows appropriately indicate the flow. Within the Bi-PAN, products are categorized based on their assembly or disassembly role as follows:
@@ IV Product Digital Twin Based on the Bi-PAN for the Electric Vehicle Battery
An EV battery serves as an exemplary case for demonstrating the PDT based on the Bi-PAN approach. Throughout the EV battery life-cycle, all three scenarios of the end-of-life phase are relevant: recycling to reintegrate scarce materials within the circular economy, repairing or replacing non-functional parts for re-use of EV batteries, and transforming EV batteries into modular units for secondary energy-storage applications. Remanufacturing or recycling of EV batteries also presents challenges, including the presence of hazardous substances and the risk of spontaneous combustion. It is therefore important to have relevant information and to select the appropriate remanufacturing/recycling route accurately and efficiently. This underscores the importance of PDT.
@@
When it comes to selecting the appropriate recycling process for an EV battery, understanding the battery structure is paramount. This entails not just knowing the list of individual components but also comprehending the interdependencies between these components, including their composition. For the creation of the PDT, the utilization of the AutomationML tool proves invaluable. This tool enables not only the listing of individual components of the EV battery but also the depiction of their dependencies and detailed information. For instance, it can provide data about each component’s position in the battery coordinate system, which proves beneficial when considering disassembly by a robot. An example of PDT and its structure in the AutomationML system unit class is depicted in Fig. 2. The data obtained from AutomationML is important, but for the ability to plan assembly or disassembly operations, we lack the linking of the already described product structure with specific processes and resources. The Bi-PAN will serve us for this purpose.
@@
Figure 2: Example of a system unit class representing the BMW i3 battery in AutomationML as a screenshot from the AutomationML Editor.
@@
As already introduced in Sec. III, Bi-PAN includes connections between products and processes, while also assigning the resources needed for each process. The Bi-PAN for the EV battery is depicted in the already mentioned Fig. 1. The following description explains the assembly process of an EV battery that starts from the top, and therefore the relevant direction for the black arrows is downwards. The assembly commences with three separate products: screws (Fastener Product), the cooling system (SubProduct), and the battery box (Elementary Product). These products are inserted into a single assembly through a process requiring resource skills for manipulation and screwing, as indicated by yellow markers. The output of this process is Stage 1 (i.e., Product in the Stage in the Bi-PAN notation) of the EV battery assembly. Stage 1 then proceeds to the next process along with bolts, BMS & cables, and eight modules, where they are once again inserted into a single assembly. This process involves resource skills manipulation, screwing, and connecting cables. The result is Stage 2 of the EV battery assembly. Stage 2 and bolts, along with five braces, proceed to the next process where they are inserted into a single unit. This process requires resource skills for manipulation and screwing. The output of this process is Stage 3 of the EV battery assembly process. Stage 3, along with the isolation blanket, enters the subsequent process, requiring resource skills for manipulation only. This creates Stage 4 of the EV battery assembly. Finally, Stage 4, along with screws and a cover, enters the last process, requiring resource skills for manipulation and screwing. The output of this process is the Final Product - the assembled EV battery.
@@
To illustrate the application of the Bi-PAN in remanufacturing, we’ve selected the scenario of replacing a broken module with a new one. This specific process represents a sequence of operations that commences in a particular state of the EV battery and concludes at the desired state. The remanufacturing process depicted in Fig. 3 is derived from the generic Bi-PAN for the entire EV battery (i.e., from Bi-PAN in Fig. 1, which is repeated without labels on the left). The flow of the whole remanufacturing process is depicted by a red arrow starting from the EV battery (Final Product) and continuing as a disassembly process to Stage 2, where the damaged module is removed and replaced by the new one, and it continues as an assembly process and ends as an EV battery (Final Product).